OpenStudy (anonymous):

If F(x) = x + 4 and G(y) =, what is the domain of G(F(x))? A. x -4 B. All real numbers. C. x 4 D. All real numbers except -4

OpenStudy (anonymous):

depends a lot on what \(G\) is

OpenStudy (anonymous):

g(y)= 1/y

OpenStudy (anonymous):

then \[g(y)=G(x+4)=\frac{1}{x+4}\]

OpenStudy (anonymous):

you cannot divide by zero (you will explode) solve \[x+4=0\] for \(x\)

OpenStudy (anonymous):

I think it's D

OpenStudy (anonymous):

It is D!! THANK U!!!

OpenStudy (anonymous):

No problem :)